# bibliography.bib

### v1

@@ 2.3 Canonical extensions and the Prime Ideal Theorem
Canonical extensions for Boolean algebras were introduced by Jónsson and Tarski (see [jonsson51i] and [jonsson51ii]) in dealing with Boolean algebras with operators. The canonical extension may be seen as a pointfree version of the lattice of compact opens into the lattice of saturated sets of a Stone space. In fact, the Prime Ideal Theorem implies that the canonical extension of a Boolean algebra $B$ is the powerset of its dual space. Without assuming the Prime Ideal Theorem, the canonical extension of a Boolean algebra may be seen as the Booleanization of $\mathcal{U}(\mathsf{Filt}(B){\setminus}\{B\})$ – or, equivalently, the Booleanization of $\mathcal{U}(\mathsf{Idl}(B))$ – as shown in [bezhanishvili22]. Canonical extensions have also been introduced for distributive lattices. On this topic, we refer the reader to [gehrke94], [gehrke01], and [gehrke08]. If we assume the Prime Ideal Theorem, the canonical extension of any distributive lattice is isomorphic to the lattice of saturated sets of the corresponding coherent space. With no choice principles this is not necessarily the case.
@@
    Compactness: for every Scott-open filter $F$ we have $\bigwedge f[F]\leq f(a)$ implies $a\in F$, for each $a\in L$.
@@ 3 Raney extensions
. We will sometimes use the expression Raney extension to refer to the coframe component of the pair, and for a pair $(L,C)$ we will say that this is a Raney extension of $L$, or that it is a Raney extension over $L$. A morphism of Raney extensions $f:(L,C)\rightarrow(M,D)$ is a coframe map $f:C\rightarrow D$ such that whenever $a\in L$ we have $f(a)\in M$ and such that the restriction $f|_{L}:L\rightarrow M$ is a frame map. We call $\mathbf{Raney}$ the category of Raney extensions with Raney maps.
@@
We shall see more concrete examples of Raney extensions in Subsection 3.1. Observe that for any Raney extension $(L,C)$, by the universal property of the ideal completion of a distributive lattice, there is a coframe surjection $\mathsf{Filt}(L)^{op}\to C$. The fact that it extends the map $L\subseteq C$, and that it preserves all meets, means that this map is defined as $F\mapsto\bigwedge F$ for each filter $F\in\mathsf{Filt}(L)$. As it is a coframe map, it has a left adjoint. Now, notice that for each filter $F\in\mathsf{Filt}(L)$ and each $c\in C$ we have:
$c\leq\bigwedge F\mbox{ if and only if }c\leq f\mbox{ for all $f\in F$, if and only if $F\subseteq{\uparrow}c\cap L$}.$
The equivalences above mean that the left adjoint we are looking for acts as $c\mapsto{\uparrow}c\cap L$, and from now on we will denote this map as ${\uparrow}^{L}:C\to\mathsf{Filt}(L)^{op}$. The fixpoints on the coframe of filters are then exactly those of the form ${\uparrow}^{L}c$ for some $c\in C$. As the starting map $\bigwedge$ is a coframe map, these form a subcolocale, which we will hereon call $C^{*}\subseteq\mathsf{Filt}(L)^{op}$. On the other hand, all elements of $C$ are fixpoints, as $\bigwedge:\mathsf{Filt}(L)^{op}\to C$ is a surjection. Let us state these observations as a theorem.
@@
For a Raney extension $(L,C)$, we have an adjunction
@@
    $\mathcal{F}$-compactness: whenever $\bigwedge f[F]\leq f(a)$ we also have $a\in F$ for every $F\in\mathcal{F}$ and every $a\in L$.
@@
For a sober space $X$, the pair $(\Omega(X),\mathcal{U}(X))$ is a $\mathcal{SO}$-canonical Raney extension, provided that the Prime Ideal Theorem holds. This is observed in Example 3.5 of [jakl20]. On the one hand, the coframe $\mathcal{U}(X)$ is join-generated by elements of the form ${\uparrow}x$ for $x\in X$, and these are intersections of neighborhood filters, which are completely prime, hence Scott-open. Then, the extension is $\mathcal{SO}$-dense. For $\mathcal{SO}$-compactness, we rely on the Hofmann-Mislove Theorem. If $F\subseteq\Omega(X)$ is a Scott-open filter, then by the Theorem it must be $\{U\in\Omega(X):\bigcap F\subseteq U\}$, and so, indeed, for every open $U$ we have $\bigcap F\subseteq U$ implies $U\in F$. Recall that the Hofmann-Mislove Theorem is dependent on the Prime Ideal Theorem, see for instance [erne18], Theorem 3. With Proposition 3.11 we will see that if we replace Scott-open by completely prime, we have an analogous result which does not rely on the Prime Ideal Theorem. We will explore the relation between the Prime Ideal Theorem and $\mathcal{SO}$-canonicity in Section 5.
Theorem 3.1 tells us that for a Raney extension $(L,C)$ we may identify elements of $C$ with filters of $L$. This proves to be useful when dealing with density and compactness. In the following, for a collection of filters $\mathcal{F}$, we denote as $\mathcal{F}^{*}$ the collection $\{{\uparrow}^{L}\bigwedge F:F\in\mathcal{F}\}$.
@@
Let us prove the first claim. If $(L,C)$ is $\mathcal{F}$-dense, then for all $c\in C$ we have that $c=\bigvee_{i}\bigwedge F_{i}$ for some collection $F_{i}\in\mathcal{F}$, that is, ${\uparrow}^{L}c={\uparrow}^{L}\bigvee_{i}\bigwedge F_{i}$. As ${\uparrow}^{L}:C\to\mathsf{Filt}(L)^{op}$ is a left adjoint, it preserves all joins, and so ${\uparrow}^{L}\bigvee_{i}\bigwedge F_{i}=\bigcap_{i}{\uparrow}^{L}\bigwedge F$. For the converse, suppose that $C^{*}\subseteq\mathcal{I}(\mathcal{F}^{*})$. We have that, for $c\in C$, ${\uparrow}^{L}c=\bigcap_{i}{\uparrow}^{L}\bigwedge F_{i}$ for some collection $F_{i}\in\mathcal{F}$. Again, by preservation of joins of ${\uparrow}^{L}$, we obtain $c=\bigvee_{i}\bigwedge F_{i}$. To see the equivalence stated in the second claim, we observe that for any filter $F\subseteq L$ we always have $F\subseteq{\uparrow}^{L}\bigwedge F$. For any collection $\mathcal{F}\subseteq\mathsf{Filt}(L)$ it is the case that for all $F\in\mathcal{F}$ we have the reverse set inclusion if and only if the Raney extension is $\mathcal{F}$-compact. But this is also equivalent to having that all filters in $\mathcal{F}$ are fixpoints of ${\uparrow}^{L}\dashv\bigwedge$, i.e. them being elements of $C^{*}$. ∎
@@
Existence and uniqueness of $\mathcal{F}$-canonical extensions of lattices to complete lattices are well-known, and these results stem from the theory of polarities by Birkhoff (see [birkhoff40]). For a general version of the existence and uniqueness results, see for instance Section 2 of [gehrke06], see [gehrke01] for its application to distributive lattices. From particularizing the analysis of [gehrke01] to the case where we start from a frame, we directly obtain the following.
@@
We now wish to adapt the theorem above to prove existence and uniqueness of $\mathcal{F}$-canonical Raney extensions on a frame $L$. An extension as required will not exist for all choices of $\mathcal{F}$, because for some choices of $\mathcal{F}$ the structure $(L,\mathcal{I}(\mathcal{F})^{op})$ will not be a Raney extension.
@@
Suppose that $(L,C)$ is a Raney extension. All the following hold.
@@
For the first item, we only notice that it is clear that $a=\bigwedge{\uparrow}^{L}a$ for all $a\in L$. The second item is one of the observations before Theorem 3.1. Let us show the third item. Suppose that $F\in C^{*}$, and that $x_{i}\in F$ is a family such that the meet $\bigwedge^{L}_{i}x_{i}$, as calculated in $L$, is strongly exact. By definition of Raney extension, this meet is preserved by the embedding $e:L\subseteq C$. This means that $\bigwedge^{L}_{i}x_{i}=\bigwedge_{i}x_{i}$, where the second meet is computed in $C$. Therefore, since $\bigwedge F\leq x_{i}$ for all $i\in I$, we also have $\bigwedge F\leq\bigwedge_{i}x_{i}$. Since $F\in C^{*}$, we have $F={\uparrow}^{L}\bigwedge F$, and so $\bigwedge^{L}_{i}x_{i}\in F$. ∎
@@
For a frame $L$ and any collection $\mathcal{F}\subseteq\mathsf{Filt}(L)$ of filters, the $\mathcal{F}$-canonical Raney extension exists if and only if:
@@
In case it exists, it is unique, up to isomorphism. Concretely, it is described as the structure coming from the theory of polarities of Birkhoff, that is, the pair $(L,\mathcal{I}(\mathcal{F})^{op})$.
@@ 3.1 Notable examples of Raney extensions
We use the characterization in Theorem 3.6. That the collections of filters below are subcolocales of $\mathsf{Filt}_{\mathcal{SE}}(L)^{op}$ follows from Theorem 2.9. The collections of exact and of strongly exact filters contain the principal filters for every frame, whereas for the last three items we refer to Proposition 2.10. ∎
Because of the isomorphisms in Theorem 2.6, for any frame $L$ the following embeddings into coframes are Raney extensions, up to isomorphism.
@@
We look at Raney extensions in relation to the concrete lattice of saturated sets of a space. We now expand a result in [jakl24] (Lemma 5.4) which states an isomorphism $\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))^{op}\cong\mathcal{U}(\mathsf{pt}(L))$, and show that it is the restriction of an adjunction which may be seen as the point-set version of that of Theorem 2.6. In the following $\varphi$ is the spatialization map of the frame $L$.

### v2

@@ 2.3 Canonical extensions and the Prime Ideal Theorem
    Compactness: for every Scott-open filter $F$, $\bigwedge f[F]\leq f(a)$ implies $a\in F$, for each $a\in L$.
@@
2.4 TD duality
A topological space $X$ is said to be $T_{D}$ if for every point $x\in X$ there are opens $U$ and $V$ such that $U{\setminus}V=\{x\}$. For a frame $L$ we say that a prime $p\in L$ is covered if whenever $\bigwedge_{i}x_{i}=p$ for some family $x_{i}\in L$ then $x_{i}=p$ for some $i\in I$. In [banaschewskitd] the $T_{D}$ spectrum of a frame $L$ is defined as the collection of covered primes of a frame, with the subspace topology inherited from the prime spectrum of $L$. This space is denoted as $\mathsf{pt}_{D}(L)$. This turns out to always be a $T_{D}$ space. A frame morphism $f:L\to M$ is a D-morphism if for every covered prime $p\in L$ the prime $f_{*}(p)$ is covered. We call $\mathbf{Frm}_{D}$ the category of frames and D-morphisms. There is a dual adjunction $\Omega:\mathbf{Top}\leftrightarrows\mathbf{Frm}_{D}:\mathsf{pt}_{D}$, where the fixpoints on the space side are the $T_{D}$ spaces, and on the frame side these are the D-spatial frames, which can be characterized as those frame such that all their elements are the meet of the covered primes above them. We will use the following two results.
Proposition 2.17.
([banaschewskitd], Proposition 2.3.2) A space $X$ is $T_{D}$ if and only if all elements of the form $X{\setminus}\overline{\{x\}}$ are covered primes in $\Omega(X)$.
Furthermore, in [arrieta21] the notion of D-sublocale is introduced. This is a sublocale $S\subseteq L$ such that the corresponding surjection is in $\mathbf{Frm}_{D}$.
Theorem 2.18.
For a frame $L$, the D-sublocales form a subcolocale $\mathsf{S}_{D}(L)\subseteq\mathsf{S}(L)$. We also have a subcolocale inclusion $\mathsf{S}_{\mathfrak{c}}(L)\subseteq\mathsf{S}_{D}(L)$.
@@ 3 Raney extensions
. We will sometimes use the expression Raney extension to refer to the coframe component of the pair, and for a pair $(L,C)$ we will say that this is a Raney extension of $L$, or that it is a Raney extension over $L$. A morphism of Raney extensions $f:(L,C)\rightarrow(M,D)$ is a coframe map $f:C\rightarrow D$ such that, whenever $a\in L$, $f(a)\in M$ and such that the restriction $f|_{L}:L\rightarrow M$ is a frame map. We call $\mathbf{Raney}$ the category of Raney extensions with Raney maps.
@@
Observe that for any Raney extension $(L,C)$, by the universal property of the ideal completion of a distributive lattice, there is a coframe surjection $\bigwedge:\mathsf{Filt}(L)^{op}\to C$ extending $L\subseteq C$. For $c\in C$, define ${\uparrow}^{L}c$ as ${\uparrow}c\cap L$. Notice that for each filter $F\in\mathsf{Filt}(L)$ and each $c\in C$:
$c\leq\bigwedge F\mbox{ if and only if }c\leq f\mbox{ for all $f\in F$, if and only if $F\subseteq{\uparrow}^{L}c$}.$
This means that ${\uparrow}^{L}:C\to\mathsf{Filt}(L)^{op}$ is left adjoint to $\bigwedge:\mathsf{Filt}(L)^{op}\to C$. As the starting map $\bigwedge$ is a coframe map, the inclusion of the fixpoints $C^{*}:=\{{\uparrow}^{L}c\mid c\in C\}\subseteq\mathsf{Filt}(L)$ is a sublocale.
@@
For a Raney extension $(L,C)$, there is an adjunction
@@
    $\mathcal{F}$-compactness: $\bigwedge f[F]\leq f(a)$ implies $a\in F$ for every $F\in\mathcal{F}$ and every $a\in L$.
@@
For a sober space $X$, the pair $(\Omega(X),\mathcal{U}(X))$ is a $\mathcal{SO}$-canonical Raney extension, provided that the Prime Ideal Theorem holds. This is observed in Example 3.5 of [jakl20]. On the one hand, the coframe $\mathcal{U}(X)$ is join-generated by elements of the form ${\uparrow}x$ for $x\in X$, and these are intersections of neighborhood filters, which are completely prime, hence Scott-open. Then, the extension is $\mathcal{SO}$-dense. For $\mathcal{SO}$-compactness, we rely on the Hofmann-Mislove Theorem. If $F\subseteq\Omega(X)$ is a Scott-open filter, then by the Theorem it must be $\{U\in\Omega(X):\bigcap F\subseteq U\}$, and so, indeed, for every open $U$, $\bigcap F\subseteq U$ implies $U\in F$. Recall that the Hofmann-Mislove Theorem is dependent on the Prime Ideal Theorem, see for instance [erne18], Theorem 3. With Proposition 5.3 we will see that if we replace Scott-open by completely prime, there is an analogous result which does not rely on the Prime Ideal Theorem. We will explore the relation between the Prime Ideal Theorem and $\mathcal{SO}$-canonicity in Subsection 7.1.
Theorem 3.1 tells us that for a Raney extension $(L,C)$ we may identify elements of $C$ with filters of $L$. In the following, for a collection of filters $\mathcal{F}$, we denote as $\mathcal{F}^{*}$ the collection $\{{\uparrow}^{L}\bigwedge F:F\in\mathcal{F}\}$.
@@
Let us prove the first claim. If $(L,C)$ is $\mathcal{F}$-dense, then, for all $c\in C$, $c=\bigvee_{i}\bigwedge F_{i}$ for some collection $F_{i}\in\mathcal{F}$, that is, ${\uparrow}^{L}c={\uparrow}^{L}\bigvee_{i}\bigwedge F_{i}$. As ${\uparrow}^{L}:C\to\mathsf{Filt}(L)^{op}$ is a left adjoint, it preserves all joins, and so ${\uparrow}^{L}\bigvee_{i}\bigwedge F_{i}=\bigcap_{i}{\uparrow}^{L}\bigwedge F$. For the converse, suppose that $C^{*}\subseteq\mathcal{I}(\mathcal{F}^{*})$. For $c\in C$, ${\uparrow}^{L}c=\bigcap_{i}{\uparrow}^{L}\bigwedge F_{i}$ for some collection $F_{i}\in\mathcal{F}$. Again, by preservation of joins of ${\uparrow}^{L}$, we obtain $c=\bigvee_{i}\bigwedge F_{i}$. To see the equivalence stated in the second claim, we observe that for any filter $F\subseteq L$ we always have $F\subseteq{\uparrow}^{L}\bigwedge F$. For any collection $\mathcal{F}\subseteq\mathsf{Filt}(L)$ it is the case that for all $F\in\mathcal{F}$ the reverse set inclusion holds if and only if the Raney extension is $\mathcal{F}$-compact. But this is also equivalent to having that all filters in $\mathcal{F}$ are fixpoints of ${\uparrow}^{L}\dashv\bigwedge$, i.e. them being elements of $C^{*}$. ∎
@@
Existence and uniqueness of what we called $\mathcal{F}$-canonical extensions of lattices to complete lattices are well-known, and these results stem from the theory of polarities by Birkhoff (see [birkhoff40]). For a general version of the existence and uniqueness results, see for instance Section 2 of [gehrke06], see [gehrke01] for its application to distributive lattices. From particularizing the analysis of [gehrke01] to the case where we start from a frame, we directly obtain the following.
@@
We now wish to adapt the theorem above to prove existence and uniqueness of $\mathcal{F}$-canonical Raney extensions on a frame $L$ for certain collections of filters $\mathcal{F}$.
@@
Suppose that $(L,C)$ is a Raney extension. Then:
@@
For the first item, we only notice that it is clear that $a=\bigwedge{\uparrow}^{L}a$ for all $a\in L$. For the second, suppose that $F\in C^{*}$, and that $x_{i}\in F$ is a family such that the meet $\bigwedge^{L}_{i}x_{i}$, as calculated in $L$, is strongly exact. By definition of Raney extension, this meet is preserved by the embedding $e:L\subseteq C$. This means that $\bigwedge^{L}_{i}x_{i}=\bigwedge_{i}x_{i}$, where the second meet is computed in $C$. Therefore, since $\bigwedge F\leq x_{i}$ for all $i\in I$, we also have $\bigwedge F\leq\bigwedge_{i}x_{i}$. Since $F\in C^{*}$, $F={\uparrow}^{L}\bigwedge F$, and so $\bigwedge^{L}_{i}x_{i}\in F$. ∎
@@
For a frame $L$ and any collection $\mathcal{F}\subseteq\mathsf{Filt}(L)$ of filters, the pair $(L,\mathcal{I}(\mathcal{F})^{op})$ is a Raney extension if and only if:
@@
In case these hold, $(L,\mathcal{I}(\mathcal{F})^{op})$ is the unique (up to isomorphism) $\mathcal{F}$-canonical Raney extension.
@@
Item (3) of Theorem 3.6 above tells us that for every Raney extension $(L,C)$ there is the upper bound $C^{*}\subseteq\mathsf{Filt}_{\mathcal{SE}}(L)^{op}$. There also is a lower bound.
Lemma 3.7.
Proof.
Proposition 3.8.
For a frame $L$ the pair $(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$ is a Raney extension, and $\mathsf{Filt}_{\mathcal{E}}(L)^{op}\subseteq C^{*}$ for all Raney extensions $(L,C)$.
Proof.
Principal filters are exact as they are closed under all meets. By Theorem 2.13, $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq\mathsf{Filt}(L)$ is a sublocale inclusion and all exact filters are strongly exact. Furthermore, by Lemma 2.11, $\mathsf{Filt}_{\mathcal{E}}(L)$ is closed under all intersections. Then, $(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$ is a Raney extension by Theorem 3.6. If $(L,C)$ is a Raney extension, the collection $C^{*}\subseteq\mathsf{Filt}(L)$ is a sublocale which contains all principal filters, by Lemma 3.5, and so $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq C^{*}$ by Lemma 3.7. ∎
Theorem 3.9.
Proof.
That every Raney extension belongs to the section $[\mathsf{Filt}_{\mathcal{E}}(L),\mathsf{Filt}_{\mathcal{SE}}(L)]$ follows from Theorem 3.6 and Proposition 3.8. Suppose that there is a sublocale $\mathcal{F}\subseteq\mathsf{Filt}_{\mathcal{SE}}(L)$ such that $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq\mathcal{F}$. By Lemma 3.7, $\mathcal{F}$ contains all principal filters, and so, by Theorem 3.6, the pair $(L,\mathcal{F}^{op})$ is a Raney extension. ∎
@@ 3.1 Notable examples of Raney extensions
We use the characterization in Theorem 3.6. That the collections of filters below are subcolocales of $\mathsf{Filt}(L)$ and that all filters in these collections are strongly exact follows from Theorem 2.13. Since principal filters are closed under all meets, they are exact and strongly exact. For the last three items we refer to Proposition 2.14. ∎
Note that all the Raney extensions above are $(L,\mathcal{I}(\mathcal{F})^{op})$ for some $\mathcal{F}\subseteq\mathsf{Filt}(L)$. This means that each extension $(L,\mathcal{I}(\mathcal{F})^{op})$ above is the unique $\mathcal{F}$-canonical one. Because of the isomorphisms in Theorem 2.8, for any frame $L$ the following embeddings into coframes are Raney extensions, up to isomorphism.
@@
4 Topological duality for Raney extensions
In this section, we show that there is an adjunction between $\mathbf{Raney}^{op}$ and $\mathbf{Top}$. For any coframe $C$, we define $\mathsf{pt}_{R}(C)$ to be the collection of its completely join-prime elements. For a Raney extension $(L,C)$, let us define the function $\varphi_{(L,C)}:C\rightarrow\mathcal{P}(\mathsf{pt}_{R}(C))$ as
$\varphi_{(L,C)}(a)=\{x\in\mathsf{pt}_{R}(C):x\leq a\}.$
It is easy to see that the following two properties hold:
  1. 1.
    $\varphi_{(L,C)}(\bigwedge_{i}a_{i})=\bigcap_{i}\varphi_{(L,C)}(a_{i})$,
  2. 2.
    $\varphi_{(L,C)}(\bigvee_{i}a_{i})=\bigcup_{i}\varphi_{(L,C)}(a_{i})$,
for each family $a_{i}\in L$. When the Raney extension $(L,C)$ is clear from the context, we will omit the subscript. By property 2, the elements of the form $\varphi_{(L,C)}(a)$ for $a\in L$ form a topology. We denote the topological space obtained by equipping the set $\mathsf{pt}_{R}(C)$ with this topology as $\mathsf{pt}_{R}(L,C)$, and we call it the spectrum of the Raney extension $(L,C)$. Since all elements of $C$ are meets of elements of $L$, from property 1 it follows that the elements of the form $\varphi_{(L,C)}(c)$ with $c\in C$ are the saturated sets of this space. Let us show functoriality of the assignment $(L,C)\mapsto\mathsf{pt}_{R}(L,C)$. Observe that the following is a pointfree version of Lemma 5.2.
Lemma 4.1.
For a Raney extension $(L,C)$, an element $x\in C$ is completely join-prime if and only if ${\uparrow}^{L}x$ is a completely prime filter.
Proof.
It is immediate that if $x\in C$ is completely join-prime then ${\uparrow}^{L}x$ is completely prime. For the converse, suppose that $x\in C$ is such that ${\uparrow}^{L}x$ is completely prime. Suppose that $x\leq\bigvee D$ for $D\subseteq C$. This means that ${\uparrow}^{L}\bigvee D\subseteq{\uparrow}^{L}x$. Observe that ${\uparrow}^{L}\bigvee D=\bigcap\{{\uparrow}^{L}d:d\in D\}$. As ${\uparrow}^{L}x$ is assumed to be completely prime, there must be some $d\in D$ such that ${\uparrow}^{L}d\subseteq{\uparrow}^{L}x$. This implies that $x\leq d$. ∎
Lemma 4.2.
For a morphism $f:(L,C)\rightarrow(M,D)$ of Raney extensions, if $x\in\mathsf{pt}_{R}(D)$ then $f^{*}(x)\in\mathsf{pt}_{R}(C)$.
Proof.
By Lemma 4.1, it suffices to show that for a morphism $f:(L,C)\rightarrow(M,D)$ of Raney extensions, if $x\in\mathsf{pt}_{R}(D)$ then ${\uparrow}^{L}f^{*}(x)$ is a completely prime filter of $L$. If $f^{*}(x)\leq\bigvee A$ for $A\subseteq L$, then as $f$ respects the frame operations of $L$, and because $f^{*}\dashv f$, $x\leq\bigvee\{f(a):a\in A\}$. Since $x$ is completely join-prime, there is some $a\in A$ such that $x\leq f(a)$, that is $f^{*}(x)\leq a$. ∎
Lemma 4.3.
The assignment $\mathsf{pt}_{R}:(L,C)\mapsto\mathsf{pt}_{R}(L,C)$ is the object part of a functor $\mathsf{pt}_{R}:\mathbf{Raney}^{op}\rightarrow\mathbf{Top}$ which acts on morphisms as $f\mapsto f^{*}$.
Proof.
That every morphism is mapped to a well-defined function between the set of points follows from Lemma 4.2. Continuity follows from the fact that the $f^{*}$-preimage of some $\varphi(a)$ for $a\in L$ is, expanding definitions,
$\{x\in\mathsf{pt}_{R}(D):f^{*}(x)\leq a\}=$
$\{x\in\mathsf{pt}_{R}(D):x\leq f(a)\}=\varphi(f(a)),$
and this set is indeed open in $\mathsf{pt}_{R}(D)$ as by definition of Raney morphism $f(a)\in M$. ∎
By Theorem 3.1, we may identify Raney extensions with collections of filters. Let us now see how to describe the spectrum under this identification.